i am just trying to write an equasion for : the sum of three consecutive odd numbers is 39. All i need is an equasion ,
Found 2 solutions by Alan3354, RAY100:
Answer by Alan3354(69443)

i am just trying to write an equasion for : the sum of three consecutive odd numbers is 39. All i need is an equasion ,
------------
The equation is:
n + n+2 + n+4 = 39
------------
That's the hard way, tho.
Get the average, = 39/3 = 13
That's the middle number
--> 11, 13, 15

Answer by RAY100(1637)

Let x,,,,,x+2,,,,,x+4,,,,be consecutive odd numbers
.
sum= (x) + (x+2) + (x+4)= 3x +6 = 39
.
3x=33
.
x=11
.
.
numbers are 11,13,15
.
.
check 11+13+15 = 39,,,,,,ok
.